OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-9, 9) and is perpendicular to -9x - 2y = 10. Please show all of your work.

OpenStudy (anonymous):

can you give me the slope that's perpendicular to the given line? we'll need that to get the equation of the line you require.

OpenStudy (anonymous):

m = -9x?

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

yes... i'm here... sorry i was afk for a bit...

OpenStudy (anonymous):

to get the slope of the given equation, -9y - 2y =10, you'll need to write this out in slope-intercept form, y = mx + b so you can identify the slope.... in other words, solve for y.....

OpenStudy (anonymous):

Sorry, I am at work and had to do something.

OpenStudy (anonymous):

'sawrite....:) that's why the internet is so awesome....

OpenStudy (anonymous):

Yea and if I can just understand this math , I will be awesome lol

OpenStudy (anonymous):

do you need help solving for y?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok... solving for y means you need to get the y variable by itself on one side of the equal sign... \(\large -9x - 2y =10 \) add 2y to both sides then simplify: \(\large -9x - 2y\color{red}{+2y} =10\color{red}{+2y} \) \(\large -9x =10+2y \) subtract 10 from both sides then simplify: \(\large -9x\color{red}{-10} =10+2y\color{red}{-10} \) \(\large -9x-10 =2y \) divide by 2 to both sides then simplify: \(\large \frac{-9}{2}x-\frac{10}{2} =y \) \(\large \frac{-9}{2}x-5 =y \) now your equation is in the form y = mx + b... the slope is \(\large m=-\frac{9}{2} \)

OpenStudy (anonymous):

now i need you to tell me the slope that is perpendicular to the line with a slope of \(\large m=-\frac{9}{2} \)

OpenStudy (anonymous):

ok let me think for a minute

OpenStudy (anonymous):

here's a hint: \(\large m_1 \cdot m_2=-1 \) IF \(\large m_1 \perp m_2 \)

OpenStudy (anonymous):

Where did the -1 come from?

OpenStudy (anonymous):

Hello?